# 2cm2cm1cm1cm

### v1

@@ 5.3 Constant Fiber Direct Integrals
For more details, we refer to [32, Section XIII.16] as well as [39] for the general theory.
@@
of equivalence classes of square-integrable $\mathcal{Y}$–valued functions with scalar product [^14] The scalar product is well-defined, by the polarization identity and the Cauchy-Schwarz inequality. See for instance [40, Section 7.3.2].
@@
$\left\|A\right\|_{\infty}\doteq\mathrm{ess-}\sup\left\{\left\|A\left(x\right)\right\|_{\mathrm{op}}:x\in\mathcal{X}\right\}<\infty\ .$
@@
If such an $A$ exists, then it is unique. Moreover, the mapping $A\mapsto D$ defined by the above equality is a $\ast$-homomorphism which is isometric. See [32, Theorem XIII.83]. The operators $A(x)\in\mathcal{B}(\mathcal{Y})$, $x\in\mathcal{X}$, are called the fibers of $D$ and we write
@@
For the reader’s convenience, we now give three essential properties of decomposable operators used in the paper, referring to [32, Theorem XIII.85 (a), (c) and (d)]. Note that $\sigma(X)$ denotes below the spectrum of any operator $X$ acting on some Hilbert space, as is usual.
@@
In other words, one gets Assertion (ii). Assertion (iii) is an elementary application of [32, Theorem XIII.83].
@@
There is a unique unitary transformation $\mathrm{U}:L^{2}(\mathcal{X},\mu)\otimes\mathcal{Y}\rightarrow L^{2}(\mathcal{X},\mathcal{Y},\mu)$ such that
@@ 5.5 Combes-Thomas Estimates
using Tonelli’s theorem. It then follows from the Riesz-Thorin theorem [48, Theorem 6.27] that, for any function $f\in\ell^{2}(\Lambda)\subseteq\ell^{1}(\Lambda)+\ell^{\infty}(\Lambda)$,
@@
and using the Neumann series [40, Lemma 4.24] for $-(T-\lambda\mathfrak{1})^{-1}B$, the element $\mathfrak{1}+(T-\lambda\mathfrak{1})^{-1}B$ is invertible with norm bounded by
@@ 5.6 Elementary Observations
Proof. [40, Proposition 2.17] already tells us that any increasing (decreasing) monotone net $(A_{i})_{i\in I}$ of self-adjoint elements in $\mathcal{B}(\mathcal{X})$ that is bounded from above (below) has a supremum (infimum) $A_{\infty}$ in $\mathcal{B}(\mathcal{X})$. The supremum (infimum) $A_{\infty}$ is itself also self-adjoint and is the weak operator limit of the net. This is proven by using the polarization identity and the Riesz representation theorem together with elementary estimates. To conclude the proof, it remains to show that $A_{\infty}\in\mathcal{B}(\mathcal{X})$ is the limit of the increasing net $(A_{i})_{i\in I}$ also in the strong operator topology. To this end, assume that $(A_{i})_{i\in I}$ is an increasing net and define the (decreasing) net $(B_{i})_{i\in I}$ of positive operators by $B_{i}\doteq A_{\infty}-A_{i}\geq 0$. By construction, this net converges in the weak operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$, which in turn implies that the net $(B_{i}^{1/2})_{i\in I}$ converges in the strong operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$. As the net $(B_{i}^{1/2})_{i\in I}$ is norm-bounded and $B_{i}=B_{i}^{1/2}B_{i}^{1/2}$, we then conclude that also $(B_{i})_{i\in I}$ converges in the strong operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$, that is, $A_{\infty}$ is the strong operator limit of the net $(A_{i})_{i\in I}$. If $(A_{i})_{i\in I}$ is a decreasing net then we consider the increasing net $(-A_{i})_{i\in I}$ to conclude that $(A_{i})_{i\in I}$ has a infimum $A_{\infty}=A_{\infty}^{\ast}\in\mathcal{B}(\mathcal{X})$, which is again the strong operator limit of the net $(A_{i})_{i\in I}$.
Acknowledgments: This work is supported by the COST Action CA18232 financed by the European Cooperation in Science and Technology (COST) and also by the Basque Government through the grant IT1615-22 and BERC 2022-2025 program and by the Ministry of Science and Innovation: PID2020-112948GB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”.

### v2

@@ 5.4 Constant Fiber Direct Integrals
For more details, we refer to [38, Section XIII.16] as well as [45] for the general theory.
@@
of equivalence classes of square-integrable $\mathcal{Y}$–valued functions with scalar product [^24] The scalar product is well-defined, by the polarization identity and the Cauchy-Schwarz inequality. See for instance [46, Section 7.3.2].
@@
$\left\|A\right\|_{\infty}\doteq\mathrm{ess-}\sup\left\{\left\|A\left(x\right)\right\|_{\mathrm{op}}:x\in\mathcal{X}\right\}<\infty\ .$ (164)
@@
If such an $A$ exists, then it is unique. Moreover, the mapping $A\mapsto D$ defined by the above equality is a $\ast$-homomorphism which is isometric. See [38, Theorem XIII.83]. The operators $A(x)\in\mathcal{B}(\mathcal{Y})$, $x\in\mathcal{X}$, are called the fibers of $D$ and we write
@@
For the reader’s convenience, we now give three essential properties of decomposable operators used in the paper, referring to [38, Theorem XIII.85 (a), (c) and (d)]. Note that $\sigma(X)$ denotes below the spectrum of any operator $X$ acting on some Hilbert space, as is usual.
@@
In other words, one gets Assertion (ii). Assertion (iii) is an elementary application of [38, Theorem XIII.83].
@@
There is a unique unitary transformation $\mathbf{V}:L^{2}(\mathcal{X},\mu)\otimes\mathcal{Y}\rightarrow L^{2}(\mathcal{X},\mathcal{Y},\mu)$ such that
@@ 5.6 Combes-Thomas Estimates
using Tonelli’s theorem. It then follows from the Riesz-Thorin theorem [94, Theorem 6.27] that, for any function $f\in\ell^{2}(\Lambda)\subseteq\ell^{1}(\Lambda)+\ell^{\infty}(\Lambda)$,
@@
and using the Neumann series [46, Lemma 4.24] for $-(T-\lambda\mathfrak{1})^{-1}B$, the element $\mathfrak{1}+(T-\lambda\mathfrak{1})^{-1}B$ is invertible with norm bounded by
@@ 5.7 Elementary Observations
Proof. [46, Proposition 2.17] already tells us that any increasing (decreasing) monotone net $(A_{i})_{i\in I}$ of self-adjoint elements in $\mathcal{B}(\mathcal{X})$ that is bounded from above (below) has a supremum (infimum) $A_{\infty}$ in $\mathcal{B}(\mathcal{X})$. The supremum (infimum) $A_{\infty}$ is itself also self-adjoint and is the weak operator limit of the net. This is proven by using the polarization identity and the Riesz representation theorem together with elementary estimates. To conclude the proof, it remains to show that $A_{\infty}\in\mathcal{B}(\mathcal{X})$ is the limit of the increasing net $(A_{i})_{i\in I}$ also in the strong operator topology. To this end, assume that $(A_{i})_{i\in I}$ is an increasing net and define the (decreasing) net $(B_{i})_{i\in I}$ of positive operators by $B_{i}\doteq A_{\infty}-A_{i}\geq 0$. By construction, this net converges in the weak operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$, which in turn implies that the net $(B_{i}^{1/2})_{i\in I}$ converges in the strong operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$. As the net $(B_{i}^{1/2})_{i\in I}$ is norm-bounded and $B_{i}=B_{i}^{1/2}B_{i}^{1/2}$, we then conclude that also $(B_{i})_{i\in I}$ converges in the strong operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$, that is, $A_{\infty}$ is the strong operator limit of the net $(A_{i})_{i\in I}$. If $(A_{i})_{i\in I}$ is a decreasing net then we consider the increasing net $(-A_{i})_{i\in I}$ to conclude that $(A_{i})_{i\in I}$ has a infimum $A_{\infty}=A_{\infty}^{\ast}\in\mathcal{B}(\mathcal{X})$, which is again the strong operator limit of the net $(A_{i})_{i\in I}$.
Acknowledgments: This work is supported by the Basque Government through the grant IT1615-22 and BERC 2022-2025 program, by the Ministry of Science and Innovation: PID2020-112948GB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, as well as by CNPq (303682/2025-6). We would also like to express our gratitude to the reviewers for their constructive suggestions, particularly the discussion of enhanced binding in Quantum Field Theory in the introduction. We are also grateful for their interest in improving the paper and for their help in correcting a calculation error by providing an explicit proof of Proposition 2.1.